OpenStudy (anonymous):

which ordered pair is a solution to the equation y=13-x?

OpenStudy (rock_mit182):

options ?

OpenStudy (anonymous):

(-1,12) (3,10) (8,5) (11, -2)

OpenStudy (inkyvoyd):

For this, you just substitute the values that you were given into the equation. The first number in the pair is x, and the second is y.

OpenStudy (inkyvoyd):

I'll work out the first answer choice for you for (-1,12) x=-1, and y=12 then the equation in question is (12) =? 13-(-1) or 12 =? 13+1 but 12 is not equal to 14, so the first ordered pair is not a solution. Can you do the second one?

OpenStudy (paki):

First things first, an ordered pair is written as (x,y)... To find a pair which satisfies this equation, plug the x value of the ordered pair into y=13-x in the x position... If the y value you get doesn't match what is given in the ordered pair, try another... In this case, the answer is (3, 10)... as y=13-10=+3 y=3 when x is 3... @ORIANAS